# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

### v1

@@ 7. Proofs of Main Results
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
For $h=\partial_{t}v,\partial_{x}v,\partial_{x}^{2}v,b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$, we again use Itô’s lemma to obtain
$h$ $(s,X^{\mathscr{G}}_{s})-h(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.5)
$=\int_{t_{i-1}}^{s}\left(\partial_{t}h(u,X^{\mathscr{G}}_{u})+b^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u$
$\quad+\int_{t_{i-1}}^{s}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}^{2}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u+\int_{t_{i-1}}^{s}(\partial_{x}h(u,X^{\mathscr{G}}_{u}))^{\top}\sigma(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\mathrm{d}W_{u}.$
By (H.3), Lemma 3.1 and due to $b,\sigma\sigma^{\top}\in C_{p}^{2,0}$, the integrands on the right hand side of the above equation have finite moments. Therefore, if we substitute (7.5) to the first and last two terms of (7.4), then every time integral term is bounded by $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$ for constant $C_{p}>0$ depends only on $b,\sigma,\bm{\pi},T,p$, and the Itô integral terms vanish. For the second and third terms of (7.4), let $h_{1}=b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$ and $h_{2}=\partial_{x}v$. We have
$\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(s,X^{\mathscr{G}}_{s})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$ (7.6)
$=$ $\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$
$+\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left(h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right].$
For the first term of (7.6), we again use the Itô’s Lemma to obtain the upper bound $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$. Note that by Hölder’s inequality and (7.5), the second term of (7.6) is bounded by
$\sqrt{\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]}\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}.$
Thus $\mathbb{E}\left[e_{i}\right]\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$ where $C_{p}$ is a constant depending only on $b,\sigma,\bm{\pi},T,p$. The desired result follows from $\sum_{i=1}^{j}(t_{i}-t_{i-1})^{2}\leq T|\mathscr{G}|$. ∎
@@
The following lemma is similar to [23, Lemma 14.1.6]. One can combine the method in Theorem 4.1 to prove it, so we omit the proof here.
@@
Suppose (H.1) holds, $l\in(0,1)\cup(1,2)\cup(2,3)$ and $g\in\mathcal{H}_{T}^{(l)}$. Then there exist a constant $C$, such that for all $s\in[0,T]$,
@@
Proof.
It suffices to prove $|\mathbb{E}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}[f(\widetilde{X}_{t^{\prime}})]|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}$ for any $t^{\prime}\in[0,T]$. Under (H. 4), by [24, Chapter 4, Theorem 5.1], for any $f\in\mathcal{H}^{(l+2)}$, the PDE (4.4) has a unique solution $v(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$. Note that $\mathcal{H}_{T}^{(l)}\subset\mathcal{H}_{s}^{(l)}$ for every $0\leq t^{\prime}\leq T$. Without loss of generality, suppose that $t^{\prime}=T$. For simplicity, we drop the notation $t^{\prime}$ and use $v$ to represent the solution of (4.4) with $t^{\prime}=T$. Recall (7.2),
@@
Recall $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$ as defined in Section 4.2. To facilitate the proof, we define $\mathcal{F}^{\xi}_{i}:=\sigma\{\xi_{j}\mid j\leq i-1\}$ for all $i\geq 0$, and define the filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with ~$\widetilde{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W}\vee\mathcal{F}^{\xi}$. Note that $\{W_{s}\}_{s\geq 0}$ is still a Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ with filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ and we have $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}\big{[}\cdot\big{|}\widetilde{\mathcal{F}}_{0}\big{]}$.
@@
As in the proof of Theorem 4.1, noting that the Itô integral in (7.5) is mean zero under $\mathbb{P}^{W}$, we can decompose $\mathbb{E}^{W}\left[e_{i}\right]$ into two terms and obtain
@@
$A_{i}:=$ $\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
@@
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
We again use the Itô’s lemma to analysis every difference term of (7.13). Note that every term in $B_{j}$ can be written as double integrals with respect to time as in the proof of Theorem 4.1, and by conditions (H.5) and (H.3), the integand is bounded by $C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{s}|^{m_{p}})$ for some $m_{p}$ and $C_{p}$ that depend on $p$. Applying Lemma 7.2, we obtain
$|B_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$
where $C_{p}$ is a constant depending on $b,\sigma,\bm{\pi},T,p$. As a result,
@@
$S_{i}:=\bigg{(}\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ $+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.15)
@@
By conditions (H.3) and (H.5), there is a constant $C_{p}$ and an integer $m_{p}$ such that $|S_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{t_{i-1}}|^{m_{p}})$. Again using Lemma 7.2, we obtain
$|A_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1}).$ (7.17)
By [35, Corollary 2.20], we have
@@
Let $\varrho=2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right)$, we obtain
$\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.19)
@@
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
$\quad\leq\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\right)$
@@
For the proof of (4.16), we can start from (7.12). Note that $\{A_{i}\}_{i}$ is still a martingale difference sequence. By (H.4), we have
$|A_{i}|\leq C\|v\|_{t_{j}}^{(l+2)}(t_{i}-t_{i-1})$ (7.20)
It suffices to analyze (7.13). Applying Lemmas 7.2 and 7.1, one can use a similar method in the proof of Theorem 4.2 to obtain
$|B_{i}|\leq C\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1}).$ (7.21)
As a result,
@@
By [35, Corollary 2.20], we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq t\bigg{)}\leq 2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right).$ (7.23)
Let $\varrho=2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right)$, we obtain
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.24)
@@
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
@@
We introduce the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. Let $\psi$ be a convex, nondecreasing function with $\psi(0)=0$. For a random variable $X(\omega):\Omega^{\xi}\rightarrow\mathbb{R}^{m},m\in\mathbb{N}_{+}$, the Orlicz norm is defined as
$\|X\|_{\psi}=\inf\left\{C>0:\mathbb{E}\left[\psi\left(\frac{|X|}{C}\right)\right]\leq 1\right\},$ (7.25)
where ||$\cdot$|| ⋅ | is the Euclid norm on $\mathbb{R}^{m}$, and the corresponding Orlicz space is $L_{\psi}$. Note that the $L_{p}$-norm $\|\cdot\|_{p}$ is the corresponding Orlicz norm with the function $\psi(x)=x^{p}$. Denote $\psi_{\alpha}=e^{x^{\alpha}}-1,\alpha>0$ (linear near the origin when $10<\alpha<$ 1 in order for $\psi_{\alpha}$ to be convex). We say $X$ is sub-exponential if and only if $X\in L_{\psi_{1}}$. It follows from $|x|^{p}/p!\leq e^{|x|}-1$ that
$\|X\|_{p}\leq\sqrt[p]{p!}\|X\|_{\psi_{1}}.$ (7.26)
Combining the above inequality and [34, Lemma 2.2.2], we derive the following lemma.
@@
For $p\geq 1$, if $\{X_{i}\}_{i=1}^{n}$ are sub-exponential random variables, then there is a constant $C_{p}$, such that
$\mathbb{E}\left[\max_{1\leq i\leq n}|X_{i}|^{p}\right]\leq C_{p}\log^{p}(1+n)\max_{i}\|X_{i}\|_{\psi_{1}}^{p}.$ (7.27)
Note that under condition (H.6), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ is globally Lipschitz with respect to the state $x$. By classical SDE results, e.g. [13, Theorem 7.9], we have the following lemma.
@@
By the Lipschitz condition of $\phi$, $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential if and only if $d_{A}(\phi(0,0,\xi),a_{0})$ is sub-exponential by the triangular inequality and the condition (1) of (H.6). we take $(t_{0},x_{0})=(0,0)$ for simplicity. Recall the notation $i(t):=\max\{i:t_{i}\leq t\}$. Note that
@@
Denote $\Delta X_{t}:=X^{\mathscr{G}}_{t}-\widetilde{X}_{t}$. We have
$\sup_{0\leq u\leq t}|\Delta X_{u}|$ $=\sup_{0\leq u\leq t}\bigg{|}\int_{0}^{u}\left(b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))-b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\right)\mathrm{d}s$ (7.31)
@@
It follows from the Lipschitz conditions and the Burkholder-Davis-Gundy inequality that for $p\geq 2$,

### v2

@@ 6. Proofs of Main Results
$+$ $\mathbb{E}\bigg[\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg]$
Based on (H.3), $\partial_{t}v,b^{\top}\partial_{x}v,\frac{1}{2}\operatorname*{Tr}(\sigma\sigma^{\top}\partial_{x}^{2}v)\in C_{p}^{2,0}$. It follows from Lemma 6.1 that
$\left|\mathbb{E}\left[e_{i}\right]\right|$ $\leq\int_{t_{i-1}}^{t_{i}}C\left(\|\partial_{t}v\|_{C_{p}^{2}}+\|b^{\top}\partial_{x}v\|_{C_{p}^{2,0}}+\frac{1}{2}\|\operatorname*{Tr}(\sigma\sigma^{\top}\partial_{x}^{2}v)\|_{C_{p}^{2,0}}\right)|\mathscr{G}|\mathrm{d}s$ (6.6)
$\leq C_{p}\|v\|_{C_{p}^{4}}|\mathscr{G}|(t_{i}-t_{i-1})$
for a constant $C_{p}$ depending only on $b,\sigma,p$. The desired result follows from (6.3), (6.6), and $\|v\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$ due to (H.3).
@@
The following lemma is similar to [25, Lemma 14.1.6]. One can combine the method in Theorem 4.1 to prove it, so we omit the proof here.
@@
Suppose (H.1) holds, $l\in(0,1)\cup(1,2)\cup(2,3)$ and $g\in\mathcal{H}_{T}^{(l)}$. Then there is a constant $C$, such that for all $s\in[0,T]$,
@@
Proof of Theorem 4.2.
It suffices to prove $|\mathbb{E}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}[f(\widetilde{X}_{t^{\prime}})]|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}$ for any $t^{\prime}\in[0,T]$. Under (H. 4), by [27, Chapter 4, Theorem 5.1], for any $f\in\mathcal{H}^{(l+2)}$, the PDE (4.4) has a unique solution $v(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$. Note that $\mathcal{H}_{T}^{(l)}\subset\mathcal{H}_{{\color[rgb]{0,0,0}t^{\prime}}}^{(l)}$ for every $0\leq t^{\prime}\leq T$. Without loss of generality, suppose that $t^{\prime}=T$. For simplicity, we drop the notation $t^{\prime}$ and use $v$ to represent the solution of (4.4) with $t^{\prime}=T$. Recall (6.3),
@@
Recall $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$ as defined in Section 4.2. To facilitate the proof, we define $\mathcal{F}^{\xi}_{i}:=\sigma\{\xi_{j}\mid j\leq i-1\}$ for all $i\geq 0$, and define the filtration $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with $\widetilde{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W}\vee\mathcal{F}^{\xi}$. Note that $\{W_{s}\}_{s\geq 0}$ is still a Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ with filtration $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ and we have $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}\big[\cdot\big|\widetilde{\mathcal{F}}_{0}\big]$.
@@
As in the proof of Theorem 4.1, we can decompose $\mathbb{E}^{W}\left[e_{i}\right]$ into two terms and obtain
@@
$A_{i}$= $\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
@@
$+$ $\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg]$
$+$ $\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg].$
As in Lemma 6.1, apply Itô’s lemma to every difference term on the right hand side of (6.13) and note that the resulting Itô’s integrals have mean zero under $\mathbb{P}^{W}$. Thus by (H.3) and (H.5), the integrands of (6.13) are bounded by $C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{s}|^{m_{p}})(t_{i}-t_{i-1})$ for some $m_{p}$ and $C_{p}$ that depend on $p$. This along with Lemma 6.3 implies $|B_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$ where $C_{p}$ is a constant depending on $b,\sigma,\bm{\pi},T,p$. As a result,
@@
$S_{i}=\bigg(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ $+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (6.15)
@@
By conditions (H.3) and (H.5), there is a constant $C_{p}$ and an integer $m_{p}$ such that $|S_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{t_{i-1}}|^{m_{p}})$. Again by Lemma 6.3, $|A_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})$. By the Azuma–Hoeffding inequality [41, Corollary 2.20], for all $t\geq 0$,
@@
Hence given $\rho\in(0,1)$, choosing $t\geq 0$ such that $\varrho=2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right)$ yields
$\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)\leq\varrho.$ (6.18)
@@
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t_{j}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t_{j}})]\right|>(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big(1+\sqrt{\ln(2/\varrho)}\big)\right)$
$\quad\leq\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\right)$
@@
To prove (4.16), we start from (6.12). For (6.13), note that under assumption (H.4), Lemma 6.3 remains valid, and the expectation in Lemma 6.2 can be replaced by $\mathbb{E}^{W}$. Based on these observations, one can use a similar method in the proof of Theorem 4.2 to obtain
$|B_{i}|\leq C\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1}),$
which implies that
@@
Note that $\{A_{i}\}_{i}$ is still a martingale difference sequence. By (H.4), we have
$|A_{i}|\leq C\|v\|_{t_{j}}^{(l+2)}(t_{i}-t_{i-1}).$
Again by the Azuma–Hoeffding inequality [41, Corollary 2.20], for all $t\geq 0$,
$\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq t\bigg)\leq 2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right).$
Hence given $\varrho\in(0,1)$, choosing $t\geq 0$ such that $\varrho=2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right)$ yields
$\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)\leq\varrho.$
@@
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t_{j}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t_{j}})]\right|>(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big(1+\sqrt{\ln(2/\varrho)}\big)\right)$
@@
We introduce the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. Define $\psi_{1}(x)\coloneqq e^{x}-1$ for all $x>0$, and for a random variable $X:\Omega^{\xi}\rightarrow\mathbb{R}^{m},m\in\mathbb{N}_{+}$, define the Orlicz norm
$\|X\|_{\psi_{1}}\coloneqq\inf\left\{C>0:\mathbb{E}\left[\psi_{1}\left(\frac{|X|}{C}\right)\right]\leq 1\right\},$ (6.20)
where $|\cdot|$ is the Euclidean norm on $\mathbb{R}^{m}$. By [40, Proposition 2.7.1], $X$ is sub-exponential if and only if $\|X\|_{\psi_{1}}<\infty$.
The following lemma follows directly from [39, Lemma 2.2.2] and the fact that $\|X\|_{p}\leq\sqrt[p]{p!}\|X\|_{\psi_{1}}$, due to the inequality that $|x|^{p}/p!\leq e^{|x|}-1$ for all $x\in{\mathbb{R}}^{m}$.
@@
There is a constant $C\geq 0$ such that for all sub-exponential random variables $\{X_{i}\}_{i=1}^{n}$,
$\big\|\max_{1\leq i\leq n}|X_{i}|\big\|_{\psi_{1}}\leq C\log(1+n)\max_{1\leq i\leq n}\|X_{i}\|_{\psi_{1}}.$ (6.21)
Thus, for $p\geq 1$, there is a constant $C_{p}\geq 0$ such that
$\mathbb{E}\left[\max_{1\leq i\leq n}|X_{i}|^{p}\right]\leq C_{p}\log^{p}(1+n)\max_{1\leq i\leq n}\|X_{i}\|_{\psi_{1}}^{p}.$ (6.22)
Note that under condition (H.6), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ is globally Lipschitz with respect to the state $x$. By classical SDE results, e.g. [14, Theorem 7.9], we have the following lemma.
@@
By (H.6), $d_{A}(\phi(0,0,\xi),a_{0})$ is sub-exponential. Recall the notation $i(t):=\max\{i:t_{i}\leq t\}$ and for all $t\in[0,T]$,
@@
For all $t\in[0,T]$, $\Delta X_{t}:=X^{\mathscr{G}}_{t}-\widetilde{X}_{t}$ satisfies
$\sup_{0\leq u\leq t}|\Delta X_{u}|$ (6.26)
$=\sup_{0\leq u\leq t}\bigg|\int_{0}^{u}\left(b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))-b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\right)\mathrm{d}s$
@@
By the Lipschitz continuity of $\sigma$ and the Burkholder-Davis-Gundy inequality, for all $p\geq 2$,